# TESSERACT: Eliminating Experimental Bias in Malware Classification across Space and Time (Extended Version)

### v1

@@ Abstract.
1. Introduction
Machine learning (ML) has become a fundamental tool in malware research within the academic security community, with its applications spanning a wide range of domains. These include Windows malware (Dahl et al., 2013; Tahan et al., 2012; Markel and Bilzor, 2014; Anderson and Roth, 2018), PDF malware (Laskov and Šrndić, 2011; Maiorca et al., 2012; Šrndić and Laskov, 2016), malicious URLs (Stringhini et al., 2013; Lee and Kim, 2012), malicious JavaScript (Rieck et al., 2010; Curtsinger et al., 2011), and Android malware (Arp et al., 2014; Mariconti et al., 2017; Grosse et al., 2017; Zhang et al., 2020). While the high performance figures associated with machine learning solutions might lead one to believe that the issue of malware classification is practically solved, this is far from true.
Malware classifiers are challenged by the dynamic nature of operational environments and attacking techniques. The constant emergence of new malware variants and families often leads to a degradation in the predictive accuracy of these classifiers over time (Mariconti et al., 2017). Ensuring temporal consistency is therefore a critical factor in evaluating the effectiveness of these classifiers. A skewed outcome can result if the experimental setup inadvertently provides the classifier with foreknowledge, leading to biased results (Allix et al., 2015; Miller et al., 2016).
These biases are not isolated instances but are widespread in different security domains. In this paper, we primarily focus on the domain of Android malware, arguing that many detection approaches, including those cited (Arp et al., 2014; Suarez-Tangil et al., 2017; Mariconti et al., 2017; Gascon et al., 2013; Zhang et al., 2014; Dash et al., 2016; Yuan et al., 2014; Grosse et al., 2017) and even encompassing our previous research, are often evaluated under conditions that do not accurately reflect real-world scenarios. Then in order to broaden our perspective, we also extend our investigation into the Windows and PDF malware domains by case study (Anderson and Roth, 2018; Šrndić and Laskov, 2016).
We identify two types of experimental bias: spatial bias and temporal bias. Spatial bias occurs when there are unrealistic assumptions about the ratio of goodware to malware in the dataset, which must be consistently enforced during the testing phase to mimic a realistic scenario. The ratio of goodware to malware varies by domain. For example, most Android apps in the wild are goodware (Lindorfer et al., 2014; Google, 2018; Allix et al., 2016b), whereas most URLs in software download events are malicious (Maggi et al., 2013; Rahbarinia et al., 2017). Temporal bias occurs when there is a temporal inconsistency between the training and testing datasets, which can lead to the integration of future knowledge into the training phase or create unrealistic settings, especially for concept drift detection (Allix et al., 2015; Miller et al., 2016; Barbero et al., 2022). This issue is particularly concerning in the case of families of closely related malware, where including even one variant in the training set may enable the algorithm to identify many variants in the testing.
@@ 1. Introduction
Prior work has investigated challenges and experimental bias in security evaluations (Rossow et al., 2012; Sommer and Paxson, 2010; van der Kouwe et al., 2018; Allix et al., 2015; Miller et al., 2016; Axelsson, 2000). The base-rate fallacy (Axelsson, 2000) describes how evaluation metrics such as True Positive Rate () and False Positive Rate () are misleading in intrusion detection, due to significant class imbalance (e.g., most traffic is benign); in contrast, we identify and address experimental settings that give misleading results regardless of the adopted metrics—even when correct metrics are reported. Previous works by Sommer and Paxson (Sommer and Paxson, 2010), Rossow et al. (Rossow et al., 2012), and Kouwe et al. (van der Kouwe et al., 2018) propose guidelines for sound security evaluations, but none of them identify temporal and spatial bias or quantify the impact of errors on classifier performance. While Allix et al. (Allix et al., 2015) and Miller et al. (Miller et al., 2016) identify initial temporal constraints in Android malware classification, we demonstrate that recent work following their guidelines (e.g., (Arp et al., 2014; Anderson and Roth, 2018)) still suffer from temporal and spatial bias. Our paper is the first to identify and address these sources of bias with novel, actionable constraints, metrics, and tool support.
@@
We identify two sources of bias that affect the evaluation of ML-based malware classifiers: temporal bias associated with incorrect train-test splits and spatial bias related to unrealistic assumptions in dataset distribution. As an extension of our previous conference paper, we experimentally verify our hypothesis on an extended Android dataset of 259,230 apps (with 10% malware) spanning over five years that, due to bias, performance can decrease up to 50% in practice in the well-known Android malware classifiers. Furthermore, we explore the impact of these biases on popular classifiers in other security domains, specifically targeting Windows PE (Anderson and Roth, 2018) and PDF malware (Šrndić and Laskov, 2016).
@@
We implement and publicly release the Python framework for our methodology, Tesseract. Accompanied by a comprehensive user guide. It demonstrates how our findings evaluate the performance-cost trade-offs of solutions designed to mitigate time decay, such as active learning (Settles, 2012) and conformal prediction (CP) (Saunders et al., 1998). Additionally, we implement the state-of-the-art CP framework-Transcendent (Barbero et al., 2022)-to be compatible with deep neural networks (DNN). This adaptation involves the implementation of a non-conformity measurement (NCM) based on the output from the SoftMax layer of a DNN model.
This article extends our work published at USENIX Security Symposium 2019 (Pendlebury et al., 2019). In this journal version, we focus on deepening and expanding the most promising aspects observed in our initial study on a larger dataset. A key enhancement is our exploration of the deep learning-based malware detection approach, previously referred to as ‘DL’ in (Pendlebury et al., 2019) and detailed in Appendix A.1. This version presents an in-depth analysis of the deep learning approach, particularly examining how it performs in comparison to the traditional ML method when applied to identical spatial and temporal bias. Furthermore, we extend our exploration to additional security domains, specifically Windows PE and PDF malware. Our findings reveal that, even in these domains where concept drift is less pronounced, applying our proposed tuning algorithm to the training data can still enhance detection capabilities.
@@
The structure of this article is as follows: In § 2, we present the background knowledge pertinent to our Android malware classification evaluation. Subsequently, in § 3, we identify and validate potential sources of experimental bias. Our methodology for conducting a spatio-temporal bias-free evaluation is delineated in § 4. In § 5, we assess the influence of these biases on the Windows PE and PDF domains by case study. § 6 demonstrates how Tesseract can be utilized to compare and assess the trade-offs associated with various budget-constrained strategies to mitigate time decay. Then, we discuss related research in § 7. Lastly, we discuss guidelines, assumptions, limitations, and availability in § 8 and § 9, and make conclusions in § 10.
2. Android Malware Classification
@@
2.1. Reference Algorithms in Android Classification
@@
In this study, we primarily focus on two popular malware detection approaches: Drebin and DeepDrebin. We opt for Drebin’s approach and its feature because it is a highly regarded technique in state-of-the-art machine learning security research (Barbero et al., 2022; Daoudi et al., 2022; Chen et al., 2023a). It utilizes a linear support vector machine (SVM) with high-dimensional binary feature vectors derived from efficient static analysis. Additionally, we incorporate DeepDrebin (Grosse et al., 2017), a deep learning approach that uses the same input features as Drebin. Deep learning models are known for their ability to discover latent feature spaces, thus potentially offering enhanced resilience against time decay (Goodfellow et al., 2016). Our previous work, published at USENIX Security 2019 (Pendlebury et al., 2019), indicated that deep learning frameworks could outperform traditional machine learning methods in terms of robustness, it becomes interesting to investigate whether these initial findings persist under varied evaluation scenarios.
@@
To ensure a robust experimental framework, we have thoroughly re-implemented both Drebin and DeepDrebin based on the detailed methodologies outlined in their foundational papers (Arp et al., 2014; Grosse et al., 2017). Successfully replicating baseline results from these studies highlights their scientific rigor and forms the foundation upon which we build our research. These methods are selected for their accessibility and the stable baselines they provide, allowing us to concentrate on the implications of our findings within the wider malware detection research field. The hyper-parameters of the re-implemented algorithms are detailed in Appendix A.1. In contrast, we have excluded MaMaDroid (Mariconti et al., 2017) from this study due to its limited efficacy, as observed in our previous analysis spanning 2014 to 2016. Our findings indicated that MaMaDroid’s performance had plateaued, making further exploration of its Markov-chain feature inefficient.
2.2. Estimating in-the-wild Malware Ratio
@@
The proportion of malware in a dataset can have a significant impact on the performance of a classifier (see § 3). In order to conduct unbiased experiments, it is crucial to use a dataset with a realistic percentage of malware to goodware. This can be achieved by down-sampling the majority class, for instance. However, it is important to note that each malware domain has its own unique ratio of malware to goodware that is typically encountered in the wild. It is therefore essential to determine whether malware is a minority, majority, or equal-sized class compared to goodware. For example, malware is the minority class in network traffic (Axelsson, 2000) and Android (Lindorfer et al., 2014), but it is the majority class in binary download events (Rahbarinia et al., 2017). Estimating the percentage of malware in the wild for a given domain can be challenging, but it can be informed by measurement papers, industry telemetry, and publicly available reports.
@@
In the Android ecosystem, there is variation in the proportion of malware among all the apps. For instance, industry telemetry by a key player suggests that around 6% of Android apps are malicious. On the other hand, AndRadar measurement study (Lindorfer et al., 2014) indicates that the ratio of Android malware in the wild is approximately 8%. The 2017 Google’s Android security report (Google, 2018) estimates the malware percentage to be between 6% and 10%, while in 2023 the AndroZoo (Allix et al., 2016b) dataset analysis reveals an incidence of 16.3% of malicious apps among the whole dataset. This data suggests that, on average, malware is the minority class in the Android domain. To stabilize the percentage of malware in our dataset, we choose to set it to 10%, which is an approximate average across different estimates, with monthly values ranging from 8% to 12%. This decision allows us to collect a dataset with statistically significant per-month malware samples. Aggressive under-sampling of the majority class would have decreased the statistical significance of the dataset, while oversampling of the goodware class would have been too resource-intensive. (see § 2.3)
@@
2.3. Android Dataset
In our research, we utilize AndroZoo (Allix et al., 2016b) as our data source, which is publicly available and contains more than 22.6 million Android apps when we write the article. Each sample from AndroZoo is labeled with a timestamp and many of them also include VirusTotal metadata results. This dataset is constantly updated through crawling from various markets, including Google Play Store and third-party markets such as Anzhi and AppChina. We selected AndroZoo dataset for our experiments because of its large size and extensive time-span, which allows us to perform realistic experiments that are aware of both space and time factors.
Defining goodware and malware. Ideally, goodware in software security refers to legitimate Android apps that do not contain any malicious behavior. Malware, on the other hand, refers to those that have been classified as containing malicious behavior. However, different anti-virus engines may have different criteria for what is malicious. Since AndroZoo’s metadata reports the number of positive anti-virus reports on VirusTotal (Google, 2004) for samples in the AndroZoo dataset, we chose for goodware and for malware, following Miller et al.’s (Miller et al., 2016) advice for a reliable ground-truth. About 9% of AndroZoo apps can be called grayware as they have when we write the article. We exclude grayware when constructing our dataset, as including gray samples could disadvantage classifiers whose features were designed with a different labeling threshold.
@@ 2.3. Android Dataset
Practical considerations, including the costs of feature extraction and storage space requirements, drive our decision to find a balance between the temporal coverage and representativeness of our dataset. We select the period from 2014 to 2018 for analysis, a timeframe frequently cited in the assessment of contemporary methodologies, as highlighted in (Chen et al., 2023b; Barbero et al., 2022; Chen et al., 2023a). This choice, we believe, provides substantial statistical significance, crucial for identifying and thoroughly evaluating the influence of experimental biases on model performance. In our examination of time decay within the Android domain, we adopt a monthly granularity and uniformly sample 259K AndroZoo apps from January 2014 to December 2018. Furthermore, we maintain an overall malware average of 10%, in line with the malware ratio around 2017 as detailed in § 2.2, while allowing a monthly variation between 8% and 12% to ensure some degree of fluctuation. This five-year span ensures a robust dataset, offering over 1,000 apps per month, with the exception of the final quarter of 2016, which experienced a lower application crawl rate.
@@
2.4. Feature Selection
We also confront the challenge of a large Drebin (Arp et al., 2014) feature space present in the dataset mentioned earlier, which encompasses over two million features across eight different aspects of Android apps, including permissions, intents, API calls, among others (Arp et al., 2014). This large feature space significantly increases computational demands and extends experiment duration, particularly problematic when analyzing DeepDrebin in detail. Owing to DeepDrebin’s architecture, such a vast number of two million input neurons leads to an exponential increase in the number of parameters, resulting in prolonged training/testing times.
To address this, we employ feature selection techniques that aim to reduce the size of the feature space while maintaining its effectiveness in discriminating between different samples. Previous research studies (Demontis et al., 2017; Roy et al., 2015; Chen et al., 2023a) have shown that utilizing only the top 10,000 features does not significantly affect Drebin classification performance. This is due to only a small number of features displaying significant discrimination and being assigned non-zero weights by the SVM learning algorithm used in Drebin classification.
@@ 2.4. Feature Selection
We reduce the Drebin feature space by selecting the top 10,000 features based on the largest absolute values in (Demontis et al., 2017). Subsequently, only the indices of these chosen features are retained for ongoing experiments. We validate the performance of both the Drebin and DeepDrebin classifiers post this feature selection. Testing results spanning from 2015 to 2018 are depicted in Figure 2. The left two plots, 1(a) and 1(b), illustrate the performance between Drebin’s full feature space and its reduced 10,000 feature space, while the 1(c) and 1(d) on the right are dedicated to DeepDrebin. It is important to note that after the initial model training with 2014 data, no further retraining occurs during the testing phase. Key metrics like the -score, Precision, and Recall for both malware and goodware classes are evaluated. The results show that in both the Drebin and DeepDrebin set, the performance in the reduced feature space remains largely consistent, aligning with findings from previous studies (Demontis et al., 2017; Roy et al., 2015; Chen et al., 2023a).
In summary, feature selection has been effective in addressing the computational and time-consuming challenges associated with large feature spaces, especially for deep learning approaches (Goodfellow et al., 2016). Our experiments demonstrate that reducing the Drebin feature space does not negate the occurrence of concept drift; in fact, it can enhance performance by eliminating irrelevant information. Therefore, we believe that utilizing feature selection is a practical way to improve the efficiency and effectiveness of malware classifiers. Thus, we decide to use the reduced Drebin feature space, with the most important 10,000 features, in the subsequent experiments.
3. Sources of Experimental Bias

### v2

@@ Abstract
Keywords Concept Drift, Experimental Bias, Malware Detection, Performance Decay
1 Introduction
Machine learning (ML) has become a fundamental tool in malware research within the academic security community, with its applications spanning a wide range of domains. These include Windows malware [18, 67, 47, 4], PDF malware [38, 45, 62], malicious URLs [64, 39], malicious JavaScript [56, 17], and Android malware [6, 46, 33, 75]. While the high performance figures associated with machine learning solutions might lead one to believe that the issue of malware classification is practically solved, this is far from true.
Malware classifiers are challenged by the dynamic nature of operational environments and attacking techniques. The constant emergence of new malware variants and families often leads to a degradation in the predictive accuracy of these classifiers over time [46]. Ensuring temporal consistency is therefore a critical factor in evaluating the effectiveness of these classifiers. A skewed outcome can result if the experimental setup inadvertently provides the classifier with foreknowledge, leading to biased results [2, 48].
These biases are not isolated instances but are widespread in different security domains. In this paper, we primarily focus on the domain of Android malware, arguing that many detection approaches, including those cited [6, 65, 46, 29, 74, 20, 73, 33] and even encompassing our previous research, are often evaluated under conditions that do not accurately reflect real-world scenarios. Then in order to broaden our perspective, we also extend our investigation into the Windows and PDF malware domains by case study [4, 62].
We identify two types of experimental bias: spatial bias and temporal bias. Spatial bias occurs when there are unrealistic assumptions about the ratio of goodware to malware in the dataset, which must be consistently enforced during the testing phase to mimic a realistic scenario. The ratio of goodware to malware varies by domain. For example, most Android apps in the wild are goodware [42, 32, 3], whereas most URLs in software download events are malicious [44, 55]. Temporal bias occurs when there is a temporal inconsistency between the training and testing datasets, which can lead to the integration of future knowledge into the training phase or create unrealistic settings, especially for concept drift detection [2, 48, 8]. This issue is particularly concerning in the case of families of closely related malware, where including even one variant in the training set may enable the algorithm to identify many variants in the testing.
@@ 1 Introduction
Prior work has investigated challenges and experimental bias in security evaluations [57, 61, 70, 2, 48, 7]. The base-rate fallacy [7] describes how evaluation metrics such as True Positive Rate () and False Positive Rate () are misleading in intrusion detection, due to significant class imbalance (e.g., most traffic is benign); in contrast, we identify and address experimental settings that give misleading results regardless of the adopted metrics—even when correct metrics are reported. Previous works by Sommer and Paxson [61], Rossow et al. [57], and Kouwe et al. [70] propose guidelines for sound security evaluations, but none of them identify temporal and spatial bias or quantify the impact of errors on classifier performance. While Allix et al. [2] and Miller et al. [48] identify initial temporal constraints in Android malware classification, we demonstrate that recent work following their guidelines (e.g., [6, 4]) still suffer from temporal and spatial bias. Our paper is the first to identify and address these sources of bias with novel, actionable constraints, metrics, and tool support.
@@
We identify two sources of bias that affect the evaluation of ML-based malware classifiers: temporal bias associated with incorrect train-test splits and spatial bias related to unrealistic assumptions in dataset distribution. As an extension of our previous conference paper, we experimentally verify our hypothesis on an extended Android dataset of 259,230 apps (with 10% malware) spanning over five years that, due to bias, performance can decrease up to 50% in practice in the well-known Android malware classifiers. Furthermore, we explore the impact of these biases on popular classifiers in other security domains, specifically targeting Windows PE [4] and PDF malware [62].
@@
We implement and publicly release the Python framework for our methodology, Tesseract. Accompanied by a comprehensive user guide. It demonstrates how our findings evaluate the performance-cost trade-offs of solutions designed to mitigate time decay, such as active learning [60] and conformal prediction (CP) [59]. Additionally, we implement the state-of-the-art CP framework-Transcendent [8]-to be compatible with deep neural networks (DNN). This adaptation involves the implementation of a non-conformity measurement (NCM) based on the output from the SoftMax layer of a DNN model.
This article extends our work published at USENIX Security Symposium 2019 [54]. In this journal version, we focus on deepening and expanding the most promising aspects observed in our initial study on a larger dataset. A key enhancement is our exploration of the deep learning-based malware detection approach, previously referred to as ‘DL’ in [54] and detailed in Appendix A.1. This version presents an in-depth analysis of the deep learning approach, particularly examining how it performs in comparison to the traditional ML method when applied to identical spatial and temporal bias. Furthermore, we extend our exploration to additional security domains, specifically Windows PE and PDF malware. Our findings reveal that, even in these domains where concept drift is less pronounced, applying our proposed tuning algorithm to the training data can still enhance detection capabilities.
@@
The structure of this article is as follows: In § 2, we present the background knowledge pertinent to our Android malware classification evaluation. Subsequently, in § 3, we identify and validate potential sources of experimental bias. Our methodology for conducting a spatio-temporal bias-free evaluation is delineated in § 4. In § 5, we assess the influence of these biases on the Windows PE and PDF domains by case study. § 6 demonstrates how Tesseract can be utilized to compare and assess the trade-offs associated with various budget-constrained strategies to mitigate time decay. Then, we discuss related research in § 7. Lastly, we discuss guidelines, assumptions, limitations, and availability in § 8 and § § 9, and make conclusions in § 10.
2 Android Malware Classification
@@
2.1 Reference Algorithms in Android Classification
@@
In this study, we primarily focus on two popular malware detection approaches: Drebin and DeepDrebin. We opt for Drebin’s approach and its feature because it is a highly regarded technique in state-of-the-art machine learning security research [8, 19, 14]. It utilizes a linear support vector machine (SVM) with high-dimensional binary feature vectors derived from efficient static analysis. Additionally, we incorporate DeepDrebin [33], a deep learning approach that uses the same input features as Drebin. Deep learning models are known for their ability to discover latent feature spaces, thus potentially offering enhanced resilience against time decay [30]. Our previous work, published at USENIX Security 2019 [54], indicated that deep learning frameworks could outperform traditional machine learning methods in terms of robustness, it becomes interesting to investigate whether these initial findings persist under varied evaluation scenarios.
@@
To ensure a robust experimental framework, we have thoroughly re-implemented both Drebin and DeepDrebin based on the detailed methodologies outlined in their foundational papers [6, 33]. Successfully replicating baseline results from these studies highlights their scientific rigor and forms the foundation upon which we build our research. These methods are selected for their accessibility and the stable baselines they provide, allowing us to concentrate on the implications of our findings within the wider malware detection research field. The hyper-parameters of the re-implemented algorithms are detailed in Appendix A.1. In contrast, we have excluded MaMaDroid [46] from this study due to its limited efficacy, as observed in our previous analysis spanning 2014 to 2016. Our findings indicated that MaMaDroid’s performance had plateaued, making further exploration of its Markov-chain feature inefficient.
2.2 Estimating in-the-wild Malware Ratio
@@
The proportion of malware in a dataset can have a significant impact on the performance of a classifier (see § 3). In order to conduct unbiased experiments, it is crucial to use a dataset with a realistic percentage of malware to goodware. This can be achieved by down-sampling the majority class, for instance. However, it is important to note that each malware domain has its own unique ratio of malware to goodware that is typically encountered in the wild. It is therefore essential to determine whether malware is a minority, majority, or equal-sized class compared to goodware. For example, malware is the minority class in network traffic [7] and Android [42], but it is the majority class in binary download events [55]. Estimating the percentage of malware in the wild for a given domain can be challenging, but it can be informed by measurement papers, industry telemetry, and publicly available reports.
@@
In the Android ecosystem, there is variation in the proportion of malware among all the apps. For instance, industry telemetry by a key player suggests that around 6% of Android apps are malicious. On the other hand, AndRadar measurement study [42] indicates that the ratio of Android malware in the wild is approximately 8%. The 2017 Google’s Android security report [32] estimates the malware percentage to be between 6% and 10%, while in 2023 the AndroZoo [3] dataset analysis reveals an incidence of 16.3% of malicious apps among the whole dataset. This data suggests that, on average, malware is the minority class in the Android domain. To stabilize the percentage of malware in our dataset, we choose to set it to 10%, which is an approximate average across different estimates, with monthly values ranging from 8% to 12%. This decision allows us to collect a dataset with statistically significant per-month malware samples. Aggressive under-sampling of the majority class would have decreased the statistical significance of the dataset, while oversampling of the goodware class would have been too resource-intensive. (see § 2.3)
@@
2.3 Android Dataset
In our research, we utilize AndroZoo [3] as our data source, which is publicly available and contains more than 22.6 million Android apps when we write the article. Each sample from AndroZoo is labeled with a timestamp and many of them also include VirusTotal metadata results. This dataset is constantly updated through crawling from various markets, including Google Play Store and third-party markets such as Anzhi and AppChina. We selected AndroZoo dataset for our experiments because of its large size and extensive time-span, which allows us to perform realistic experiments that are aware of both space and time factors.
Defining goodware and malware. Ideally, goodware in software security refers to legitimate Android apps that do not contain any malicious behavior. Malware, on the other hand, refers to those that have been classified as containing malicious behavior. However, different anti-virus engines may have different criteria for what is malicious. Since AndroZoo’s metadata reports the number of positive anti-virus reports on VirusTotal [31] for samples in the AndroZoo dataset, we chose for goodware and for malware, following Miller et al.’s [48] advice for a reliable ground-truth. About 9% of AndroZoo apps can be called grayware as they have when we write the article. We exclude grayware when constructing our dataset, as including gray samples could disadvantage classifiers whose features were designed with a different labeling threshold.
@@ 2.3 Android Dataset
Practical considerations, including the costs of feature extraction and storage space requirements, drive our decision to find a balance between the temporal coverage and representativeness of our dataset. We select the period from 2014 to 2018 for analysis, a timeframe frequently cited in the assessment of contemporary methodologies, as highlighted in [15, 8, 14]. This choice, we believe, provides substantial statistical significance, crucial for identifying and thoroughly evaluating the influence of experimental biases on model performance. In our examination of time decay within the Android domain, we adopt a monthly granularity and uniformly sample 259K AndroZoo apps from January 2014 to December 2018. Furthermore, we maintain an overall malware average of 10%, in line with the malware ratio around 2017 as detailed in § 2.2, while allowing a monthly variation between 8% and 12% to ensure some degree of fluctuation. This five-year span ensures a robust dataset, offering over 1,000 apps per month, with the exception of the final quarter of 2016, which experienced a lower application crawl rate.
@@
2.4 Feature Selection
We also confront the challenge of a large Drebin [6] feature space present in the dataset mentioned earlier, which encompasses over two million features across eight different aspects of Android apps, including permissions, intents, API calls, among others [6]. This large feature space significantly increases computational demands and extends experiment duration, particularly problematic when analyzing DeepDrebin in detail. Owing to DeepDrebin’s architecture, such a vast number of two million input neurons leads to an exponential increase in the number of parameters, resulting in prolonged training/testing times.
To address this, we employ feature selection techniques that aim to reduce the size of the feature space while maintaining its effectiveness in discriminating between different samples. Previous research studies [22, 58, 14] have shown that utilizing only the top 10,000 features does not significantly affect Drebin classification performance. This is due to only a small number of features displaying significant discrimination and being assigned non-zero weights by the SVM learning algorithm used in Drebin classification.
@@ 2.4 Feature Selection
We reduce the Drebin feature space by selecting the top 10,000 features based on the largest absolute values in [22]. Subsequently, only the indices of these chosen features are retained for ongoing experiments. We validate the performance of both the Drebin and DeepDrebin classifiers post this feature selection. Testing results spanning from 2015 to 2018 are depicted in Figure 2. The left two plots, 2(a) and 2(b), illustrate the performance between Drebin’s full feature space and its reduced 10,000 feature space, while the 2(c) and 2(d) on the right are dedicated to DeepDrebin. It is important to note that after the initial model training with 2014 data, no further retraining occurs during the testing phase. Key metrics like the -score, Precision, and Recall for both malware and goodware classes are evaluated. The results show that in both the Drebin and DeepDrebin set, the performance in the reduced feature space remains largely consistent, aligning with findings from previous studies [22, 58, 14].
In summary, feature selection has been effective in addressing the computational and time-consuming challenges associated with large feature spaces, especially for deep learning approaches [30]. Our experiments demonstrate that reducing the Drebin feature space does not negate the occurrence of concept drift; in fact, it can enhance performance by eliminating irrelevant information. Therefore, we believe that utilizing feature selection is a practical way to improve the efficiency and effectiveness of malware classifiers. Thus, we decide to use the reduced Drebin feature space, with the most important 10,000 features, in the subsequent experiments.
3 Sources of Experimental Bias